OpenStudy (anonymous):

To find the height of a tall tree, a surveyor moves 140 feet away from the base of the tree and then, with a transit 4 feet tall, measures the angle of elevation to the top of the tree to be 53°. What is the height of the tree?

OpenStudy (anonymous):

In order to solve this question, you need to use trigonometry and the rules of similar triangles. To find 'y', we use trigonometry. \[\tan 37 = y \div 4\] y = 4 tan37 y = 3.014 Now you need to find 'x' and that will help you find the height of the tree Using similar triangles, we get the following relation: \[(140+y)/(x+4) = 140/x\] we know y to be 3.014 so this becomes \[(143.014)/(x+4) = 140/x\] Now you cross multiply and solve for 'x' 140(x+4) = 143.014(x) 3.014x = 560 x = 185.8 feet Height of tree according to figure = x+4 height of tree is approximately 190 feet